Elas table pens What is the price elasticity of demand for pens in the price range R19 to R17?  

1
Expert's answer
2021-07-27T12:58:01-0400

Price elasticity of demand tends to be %change in quantity demanded of service or good divided by % price change.

"Pe = \frac{\frac{(Qd2-Qd1)}{Qd1}}{\frac{(P2-P1)}{P1}}"